OpenStudy (gabylovesyou):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (gabylovesyou):

@kausarsalley

OpenStudy (gabylovesyou):

@waterineyes

OpenStudy (anonymous):

I can help for first part only...

OpenStudy (gabylovesyou):

ok... how do i solve them ?

OpenStudy (anonymous):

\(x^2- 25\) = 0 Here you must know: \[a^2 - b^2 = (a + b)(a-b)\] So, above equation can be written as: \[x^2 - 5^2 = 0\] Now use that formula here..

OpenStudy (gabylovesyou):

(x + 5)(x - 5) = 0

OpenStudy (gabylovesyou):

@waterineyes

OpenStudy (jim766):

that is factored, but not solved... x + 5 = 0 or x - 5 = 0 x = -5 x = 5

OpenStudy (anonymous):

Yep, so you got x = 5 and x = -5..

OpenStudy (gabylovesyou):

ok and for the other one?

OpenStudy (anonymous):

Now, in later case: \(x^2 - 2x - 15\) = 0 Right??

OpenStudy (anonymous):

Can you make factors of 15 which will add to give -2 and product to give -15 ??

OpenStudy (gabylovesyou):

x^2 = 2x + 15

OpenStudy (gabylovesyou):

how did u get x^2 - 2x - 15 = 0 ?

OpenStudy (anonymous):

Subtract 2x from both the sides.. Then Subtract 15 from both the sides..

OpenStudy (anonymous):

Getting??

OpenStudy (gabylovesyou):

oo ok

OpenStudy (anonymous):

So, can you find what I have said??

OpenStudy (gabylovesyou):

3 and -5

OpenStudy (anonymous):

Yep: So you will get: \[(x + 3)(x - 5) = 0\] Find x here..

OpenStudy (gabylovesyou):

x = -3, x = 5

OpenStudy (anonymous):

Yep.. So these are the solutions for x and answer for first part.. And I am weak at graphing part..

OpenStudy (gabylovesyou):

what about Part 2? those answers represent what?

OpenStudy (anonymous):

Those represent that the parabola cuts x-axis at these positions, that are -5 and 5 for earlier case and 5 and -3 for later part..

OpenStudy (gabylovesyou):

im gonna send u a PM @waterineyes

OpenStudy (gabylovesyou):

sent it..